OpenStudy (anonymous):

If I buy two apples and six oranges, I will spend $5.60. However, if I buy six apples and two oranges, I will spend $4.80. How much does one orange cost? $0.55 $0.70 $0.75 $0.60

OpenStudy (anonymous):

$.75 :)

OpenStudy (anonymous):

$0.75 (EACH ORANGE) *6oranges= $4.50 $5.60-$4.50= $1.10 $1.10/2 apples= $0.55 EACH APPLE $0.55*6apples= $3.30 $4.80-$3.30= $1.50 (MONEY LEFT) $1.50/2oranges= $0.75 ANSWER: C, $0.75

OpenStudy (anonymous):

@EmmaH medal?

OpenStudy (anonymous):

2(a) + 6(o) = 5.6 6(a) + 2(o) = 4.8 Multiply the first equation by -3 -6(a) - 18(o) = -16.8 6(a) + 2(o) = 4.8 Add the first equation to the second: -16(o) = -12 Solve for "o": o = 0.75

OpenStudy (anonymous):

That very last step is gotten by dividing both sides by "-16".

OpenStudy (anonymous):

ok:) thanks

OpenStudy (anonymous):

How come I didn't get the medal?... I answered first AND showed work.. :/